# Integrating One-Shot View Planning with a Single Next-Best View via Long-Tail Multiview Sampling

### v2

@@ Abstract
Figure 1: Comparative results of reconstructing an unknown object: Each method is depicted through reconstructed 3D models (red point clouds), local paths (cyan), global paths (purple), views (red-green-blue coordinate systems), and the same initial view (black circle). In (a) and (b), it is observed that both the iterative NBV method and the one-shot set-covering view-planning (SCVP) network fail to provide complete surface details (yellow voxels in enlarged areas). To address these missing surfaces, we introduce a novel combined pipeline that incorporates four NBVs before activating the SCVP network, as shown in (c). However, this approach requires more NBVs and paths (4 NBVs with 4 long local paths). Therefore, we propose the multiview-activated MA-SCVP network, trained on our innovative long-tail multiview dataset as shown in (d), to enhance the reconstruction efficiency (requiring 1 NBV with 1 local path) while maintaining high-quality reconstruction.
footnotetext: This work was supported by the NSFC Project under Grant 61771146 (Corresponding author: Hui Wei).footnotetext: Sicong Pan, Hao Hu, and Hui Wei are with the Laboratory of Algorithms for Cognitive Models, School of Computer Science, Fudan University, China (email: span@uni-bonn.de, hao.hu@intel.com, weihui@fudan.edu.cn).footnotetext: Sicong Pan, Nils Dengler, Tobias Zaenker, Murad Dawood, and Maren Bennewitz are with the Humanoid Robots Lab, Computer Science Institute VI, University of Bonn, Germany (email: span@uni-bonn.de, dengler@cs.uni-bonn.de, tzaenker@cs.uni-bonn.de, dawood@cs.uni-bonn.de, maren@cs.uni-bonn.de).footnotetext: Hao Hu is with Intel Asia-Pacific Research & Development Ltd.
@@ I Introduction
In simulated reconstruction tests, the MA-SCVP network trained on our long-tail dataset exhibits superior overall performance compared to traditional sampling methods [8, 10]. Notably, training on our long-tail dataset also improves the surface coverage of the first few NBVs obtained from NBV networks. This enhancement facilitates a more seamless integration of the MA-SCVP with NBV networks. Ablation studies conducted on our combined pipeline reveal that, with the assistance of training on our long-tail multiview dataset, only a single NBV is required to provide additional information before activating the MA-SCVP network. Fig. 1(d) illustrates an example of the integration of one-shot view planning with a single NBV through our long-tail multiview sampling, reducing the required NBVs and paths.
@@
  • We propose a novel combined pipeline that incorporates a few NBVs before embarking on one-shot view planning. Ablation studies confirm that a single NBV is adequate for providing extra information during online data collection. Our system concurrently achieves both high-quality and high-efficiency unknown object reconstruction.
  • We deploy our system on a robotic arm equipped with an RGB-D camera in a real-world environment. Comparative experiments demonstrate the robust generalization capabilities and reconstruction performance of our system.
@@ V-A1 Object Datasets
V-A2 Test Cases
The view planning performance can be slightly sensitive to the object rotations and initial views. We randomly select five different object rotations and twenty initial views for each object model. In total, we have $30\times 5\times 20=3,000$ test cases. We set the same number of test cases for all methods to make the reconstruction comparable.
@@ V-A3 Evaluation
This work focuses on the reconstruction quality and reconstruction efficiency. The reconstruction quality is quantified by the visible surface coverage VSC:
@@
Note that the number of local paths is $RV-1$.
Furthermore, we assess the reconstruction time RT that is the total execution time. In the simulation, we conduct a more comprehensive analysis, breaking down the time into inference time and OctoMap update time.
@@ V-B3 Metrics and Results
  Iterations   PCNBV citezeng2020pc   NBVNet [10]
@@
TABLE II: Evaluation of NBV networks trained on datasets using different sampling methods: Each value is the mean VSC value with its standard deviation, computed over 3,000 tests. VSC stands for visible surface coverage. As can be seen, NBV networks trained on the Long-Tail dataset achieve higher coverage in early iterations.
@@ V-D Reconstruction Study on Combined Pipeline
TABLE IV: Evaluation of different NBV methods to identify the most effective approach in early iterations: Each value is the mean VSC value with its standard deviation, computed over 3,000 tests. VSC stands for visible surface coverage. As can be seen, PCNBV exhibits superior performance in the early iterations (1-5), making it our final choice.
@@
TABLE V: Evaluation of combined pipelines on SCVP trained on the singleview dataset [6] and MA-SCVP trained on the long-tail multiview dataset: Each value is the mean value and its standard deviation, computed over 3,000 tests. RV, VSC, and MC stand for the number of required views, visible surface coverage, and movement cost, respectively. Note that $k=0$ means the input of only the initial view (singleview-activated) while $k\geq 1$ means the input of the initial view and the additional $k$ NBVs from PCNBV (multiview-activated). As can be seen, MA-SCVP achieves a consistently high VSC with only a single NBV ($k=1$), exhibiting a stable plateau across $k=1,2,3,4,5$ NBVs combined. This makes 1-NBV+MA-SCVP our final choice and $t$-tests shown in Table VI are analyzed to further confirm our choice.
@@ V-D2 Determination of One-Shot Submodule and Number k
  • When incorporating NBVs into MA-SCVP, a single NBV ($k=1$) proves adequate to achieve a high VSC. Adding more NBVs ($k=2,3,4,5$) yields no significant benefit for VSC, as a stable plateau is observed. This trend is confirmed by the results of the $t$-tests on these pipelines presented in Table VI.
  • Furthermore, incorporating more NBVs into MA-SCVP consistently results in a substantial increase in MC. For instance, achieving a non-significant VSC improvement would necessitate an additional 50% movement (comparing $k=4$ to $k=1$), which is not deemed worthwhile.
  Method   Reference Method   Significance
RV VSC MC
    0-NBV+MA-SCVP 0-NBV+SCVP   TRUE TRUE TRUE
  1-NBV+MA-SCVP   0-NBV+MA-SCVP   FALSE TRUE TRUE
  2-NBV+MA-SCVP   1-NBV+MA-SCVP   FALSE FALSE TRUE
  3-NBV+MA-SCVP   1-NBV+MA-SCVP   FALSE FALSE TRUE
  4-NBV+MA-SCVP   1-NBV+MA-SCVP   FALSE FALSE TRUE
  5-NBV+MA-SCVP   1-NBV+MA-SCVP   TRUE FALSE TRUE
TABLE VI: $t$-tests on different combined pipelines in Table V: Each value is whether the difference between two methods is significant or not according to the student $t$-test with a $p$-value of 0.001. RV, VSC, and MC stand for the number of required views, visible surface coverage, and movement cost, respectively. As can be seen, (1) 0-NBV+MA-SCVP exhibits a significant VSC increase compared to 0-NBV+SCVP, indicating the effectiveness of our long-tail multiview sampling; (2) 1-NBV+MA-SCVP exhibits a significant VSC increase compared to 0-NBV+MA-SCVP, indicating the effectiveness of our combined pipeline with a single NBV; (3) the rest $k$-NBV+MA-SCVP ($k=2,3,4,5$) do not exhibit a significant VSC increase compared to 1-NBV+MA-SCVP but require a significant higher MC. These confirm a single NBV is sufficient for our MA-SCVP to achieve both high-quality and high-efficiency reconstruction of unknown objects.
@@
    Combined (Ours) Auto-Stop   1-NBV+MASCVP   13.01 ± 1.46 $\mathbb{{}^{\star}}$98.41 ± 2.67 3.07 ± 0.24
TABLE VII: Evaluation of three types of view planning systems: iterative NBV, one-shot, and combined. Each value is the mean value and its standard deviation, computed over 3,000 tests. $m$-based stop criterion [82, 83] finishes the reconstruction when the variation in the number of frontier voxels becomes smaller than a threshold for consecutive $m=3$ sensing operations. The threshold is set to 0.1% (strict) or 0.5% (lenient) of the total voxels. +Mov stands for the movement-cost weighted NBV evaluation. RV, VSC, and MC stand for the number of required views, visible surface coverage, and movement cost, respectively. The star indicates significance according to the student t-test with a p-value of 0.001. As can be seen, our 1-NBV+MASCVP achieves a significantly higher VSC than all baselines, while reducing MC by 45% compared to the second-highest VSC method (GMC with $<0.1\%$).
Based on these findings, we settle on the following choices: (1) PCNBV for the NBV submodule, (2) MA-SCVP for the one-shot submodule, and (3) opting for $k=1$ as the number of NBV combined. Therefore, the following 1-NBV+MA-SCVP stands for this determined pipeline.
@@ V-E3 Inference and Update Time
Table VIII reports the time required for inferring view planning once using different methods. As a learning-based method, our system achieves a short inference time. Another factor influencing the reconstruction time is the OctoMap update time, which is measured at 0.163 ± 0.111 seconds for a single update. OctoMap-based NBV methods need to be updated at each iteration (about 7 or 11 in total), while our system requires only two updates (one for the initial view and one for NBV). This also indicates that the remaining number of required views does not reduce the efficiency of our system. Overall, our system exhibits high computational efficiency, totaling around 0.564 seconds.
@@ V-F5 Comparison of System Performance
  SCVP [6]   10.2 ± 0.95 95.88 ± 4.28 1.28 ± 0.15 82.2 ± 11.7
  Ours   12.93 ± 1.27 $\mathbb{{}^{\star}}$99.35 ± 0.29 1.73 ± 0.22 105.3 ± 12.6
TABLE X: Evaluation of view planning systems with the dynamic strategy in the real world: GMC and MCMF+Mov adopt the $m$-based stop criterion ($m=3$) with a threshold 0.4%. Each value is the mean value and its standard deviation, computed over 60 tests. RV, VSC, MC, and RT stand for the number of required views, visible surface coverage, movement cost, and reconstruction time, respectively. The star indicates significance according to the student t-test with a p-value of 0.001. As can be seen, our system achieves a significantly higher VSC than baselines, while also exhibiting substantial reductions in MC compared to the second-highest VSC method (GMC).
@@
  Ours   $\mathbb{{}^{\star}}$99.35 ± 0.29 1.73 ± 0.22 105.3 ± 12.6
TABLE XI: Evaluation of applying our stop criterion to NBV systems in the real world: Each value is the mean value and its standard deviation, computed over 60 tests. RV (the number of required views) of NBV systems is 12.93 ± 1.27, which is the same as our system. VSC, MC, and RT stand for visible surface coverage, movement cost, and reconstruction time, respectively. The star indicates significance according to the student t-test with a p-value of 0.001. As can be seen, while our stop criterion markedly enhances NBV systems by increasing VSC and reducing standard deviation, our system still achieves a significantly higher VSC.
@@ B-C Stop Criterion Analysis
  1-NBV+MASCVP (Ours)   13.01 ± 1.46 $\mathbb{{}^{\star}}$98.41 ± 2.67 3.07 ± 0.24
TABLE XVI: Evaluation of random methods with a fixed number of views and our stop criterion: Each value is the mean value and its standard deviation, computed over 3,000 tests. RV, VSC, and MC stand for the number of required views, visible surface coverage, and movement cost, respectively. The star indicates significance according to the student t-test with a p-value of 0.001. As can be seen, our 1-NBV+MASCVP achieves a significantly higher VSC against random methods, indicating our effective stop criterion.

### v3

@@ Abstract
Figure 1: Comparative results of reconstructing an unknown object: Each method is depicted through reconstructed 3D models (red point clouds), local paths (cyan), global paths (purple), views (red-green-blue coordinate systems), and the same initial view (black circle). In (a) and (b), it is observed that both the iterative NBV method and the one-shot set-covering view-planning (SCVP) network fail to provide complete surface details (yellow voxels in enlarged areas). To address these missing surfaces, we introduce a novel combined pipeline that incorporates four NBVs before activating the SCVP network, as shown in (c). However, this approach requires more NBVs and paths (4 NBVs with 4 long local paths). Therefore, we propose the multiview-activated MA-SCVP network, trained on our innovative long-tail multiview dataset as shown in (d), requiring only 1 NBV with 1 local path. Our novel pipeline preserves high-quality reconstruction while reducing the number of necessary NBVs and minimizing path length.
footnotetext: This work was supported by the NSFC Project under Grant 61771146 (Corresponding author: Hui Wei).footnotetext: Sicong Pan, Hao Hu, and Hui Wei are with the Laboratory of Algorithms for Cognitive Models, School of Computer Science, Fudan University, China (email: 18210240033@fudan.edu.cn, 20210240286@fudan.edu.cn, weihui@fudan.edu.cn).footnotetext: Sicong Pan, Nils Dengler, Tobias Zaenker, Murad Dawood, and Maren Bennewitz are with the Humanoid Robots Lab, University of Bonn, Germany (email: span@uni-bonn.de, dengler@cs.uni-bonn.de, tzaenker@cs.uni-bonn.de, dawood@cs.uni-bonn.de, maren@cs.uni-bonn.de).
@@ I Introduction
In simulated reconstruction tests, the MA-SCVP network trained on our long-tail dataset exhibits superior overall performance compared to traditional sampling methods [8, 10]. Notably, training on our long-tail dataset also improves the surface coverage of the first few NBVs obtained from NBV networks. This enhancement facilitates a more seamless integration of the MA-SCVP with NBV networks. Ablation studies conducted on our combined pipeline reveal that, with the assistance of training on our long-tail multiview dataset, a single NBV is required to provide additional information before activating the MA-SCVP network. Fig. 1(d) illustrates an example of the integration of one-shot view planning with a single NBV through our long-tail multiview sampling, reducing the required NBVs and paths while maintaining high-quality reconstruction.
@@
  • We propose a novel combined pipeline that incorporates a few NBVs before embarking on one-shot view planning. Ablation studies confirm that a single NBV is adequate for providing extra information during online data collection. This helps us reduce the onboard resource demands for the required NBVs and paths.
  • We deploy our system on a robotic arm equipped with an RGB-D camera in a real-world environment. Comparative experiments demonstrate the robust generalization capabilities and reconstruction performance of our system. Our system concurrently achieves both high-quality and high-efficiency unknown object reconstruction.
@@ V-A1 Object Datasets
V-A2 Test Cases and Statistical Significance
The view planning performance can be slightly sensitive to the object rotations and initial views. We randomly select five different object rotations and twenty initial views for each object model. In total, we have $30\times 5\times 20=3,000$ test cases. Importantly, all methods were evaluated on the same test cases, allowing for two-tailed statistical significance analysis using both the paired t-test and the Wilcoxon signed-rank test (due to some non-normally distributed data), with an alpha value of 0.05.
@@ V-A3 Evaluation
The reconstruction quality is quantified by the visible surface coverage VSC:
@@
Note that the number of local paths is $RV-1$. We assess the reconstruction time RT that is the total execution time.
@@ V-B3 Metrics and Results
    Iterations     PCNBV [8]     NBVNet [10]
@@
TABLE II: Evaluation of NBV networks trained on datasets using different sampling methods: Each value is the mean VSC value with its standard deviation, computed over 3,000 tests. VSC stands for visible surface coverage. The initial VSC (iteration 0) is the same for all methods, with a value of 38.16 ± 7.55. As can be seen, NBV networks trained on the Long-Tail dataset achieve higher coverage in early iterations.
@@ V-D Reconstruction Study on Combined Pipeline
TABLE IV: Evaluation of different NBV methods to identify the most effective approach in early iterations: Each value is the mean VSC value with its standard deviation, computed over 3,000 tests. VSC stands for visible surface coverage. The initial VSC (iteration 0) is the same for all methods, with a value of 38.16 ± 7.55. As can be seen, PCNBV exhibits superior performance in the early iterations (1-5), making it our final choice.
@@
TABLE V: Evaluation of combined pipelines on SCVP trained on the singleview dataset [6] and MA-SCVP trained on the long-tail multiview dataset: Each value is the mean value and its standard deviation, computed over 3,000 tests. RV, VSC, and MC stand for the number of required views, visible surface coverage, and movement cost, respectively. Note that $k=0$ means the input of only the initial view (singleview-activated) while $k\geq 1$ means the input of the initial view and the additional $k$ NBVs from PCNBV (multiview-activated). As can be seen, MA-SCVP achieves a sufficient VSC with only a single NBV ($k=1$), exhibiting a plateau across $k=1,2,3$ NBVs combined as confirmed in Table VI. Furthermore, although adding more NBVs ($k=4,5$) improves VSC, it also substantially increases the MC. To achieve a better trade-off, we selected 1-NBV+MA-SCVP as our final choice.
@@ V-D2 Determination of One-Shot Submodule and Number k
  • When incorporating NBVs into MA-SCVP, a single NBV ($k=1$) proves adequate to achieve a high VSC. Adding a few more NBVs ($k=2,3$) yields no significant benefit for VSC, as a stable plateau is observed. This trend is confirmed by significance tests presented in Table VI.
  • While incorporating a larger number of NBVs ($k=4,5$) into MA-SCVP leads to a significant increase in VSC, it also results in a substantial increase in MC (58% and 73%). This is logical, as adding more NBVs introduces additional local paths and more online information (pure NBVs can achieve a very high VSC as in Table IV).
    Method     Reference Method     Significance of VSC     Growth Rate of MC
            t-test     Wilcoxon
[…]
    2-NBV+MA-SCVP     1-NBV+MA-SCVP     n.s.     n.s.     20%
    3-NBV+MA-SCVP     1-NBV+MA-SCVP     n.s.     n.s.     40%
[…]
@@
      Combined (Ours)     Auto-Stop     1-NBV+MASCVP     13.01 ± 1.46 ⋆⋆⋆98.41 ± 2.67 3.07 ± 0.24
TABLE VII: Evaluation of three types of view planning systems: iterative NBV, one-shot, and combined. Each value is the mean value and its standard deviation, computed over 3,000 tests. $m$-based stop criterion [82, 83] finishes the reconstruction when the variation in the number of frontier voxels becomes smaller than a threshold for consecutive $m=3$ sensing operations. The threshold is set to 0.1% (strict) or 0.5% (lenient) of the total voxels. +Mov stands for the movement-cost weighted NBV evaluation. RV, VSC, and MC stand for the number of required views, visible surface coverage, and movement cost, respectively. ⋆⋆⋆ indicates a consistent significance at $p<0.001$ between our method and all baselines (i.e., several paired tests), as determined by both paired t-test and the Wilcoxon signed-rank test. As can be seen, our 1-NBV+MASCVP achieves a significantly higher VSC than all baselines, while reducing MC by 45% compared to the second-highest VSC method (GMC with $<0.1\%$).
Based on these findings, we consider the trade-off between achieving a high VSC and using fewer NBVs and less MC, opting for MA-SCVP with $k=1$ as the number of combined NBVs for the one-shot submodule. It is worth mentioning that 1-NBV+MA-SCVP demonstrates an even higher VSC (98.41) than 5-NBV+SCVP (98.27), accompanied by a substantial MC reduction of 41%. This supports our claim that our long-tail multiview sampling effectively reduces the required NBVs and MC. The following 1-NBV+MA-SCVP stands for this determined pipeline.
@@ V-E3 Inference and Update Time
To have a more comprehensive analysis, we break down the reconstruction time into inference time and OctoMap update time. Table VIII reports the time required for inferring view planning once. As a learning-based method, our system achieves a short inference time. The OctoMap update time is measured at 0.163 ± 0.111 seconds for a single update. OctoMap-based NBV methods need to be updated at each iteration (about 7 or 11 in total), while our system requires only two updates (one for the initial view and one for NBV). This also indicates that the remaining number of required views does not reduce the efficiency of our system. Overall, our system exhibits high computational efficiency, totaling around 0.564 seconds.
@@ V-F5 Comparison of System Performance
    SCVP [6]     10.20 ± 0.95 95.88 ± 4.28 1.28 ± 0.15 82.2 ± 11.7
    Ours     12.93 ± 1.27 ⋆⋆⋆99.35 ± 0.29 1.73 ± 0.22 105.3 ± 12.6
TABLE X: Evaluation of view planning systems with the dynamic strategy in the real world: GMC and MCMF+Mov adopt the $m$-based stop criterion ($m=3$) with a threshold 0.4%. Each value is the mean value and its standard deviation, computed over 60 tests. RV, VSC, MC, and RT stand for the number of required views, visible surface coverage, movement cost, and reconstruction time, respectively. ⋆⋆⋆ indicates a consistent significance at $p<0.001$ between our method and all baselines (i.e., several paired tests), as determined by both paired t-test and the Wilcoxon signed-rank test. As can be seen, our system achieves a significantly higher VSC than baselines, while also exhibiting substantial reductions in MC compared to the second-highest VSC method (GMC).
@@
    Ours     ⋆⋆⋆99.35 ± 0.29 1.73 ± 0.22 105.3 ± 12.6
TABLE XI: Evaluation of applying our stop criterion to NBV systems in the real world: Each value is the mean value and its standard deviation, computed over 60 tests. RV (the number of required views) of NBV systems is 12.93 ± 1.27, which is the same as our system. VSC, MC, and RT stand for visible surface coverage, movement cost, and reconstruction time, respectively. ⋆⋆⋆ indicates a consistent significance at $p<0.001$ between our method and all baselines (i.e., several paired tests), as determined by both paired t-test and the Wilcoxon signed-rank test. As can be seen, while our stop criterion markedly enhances NBV systems by increasing VSC and reducing standard deviation, our system still achieves a significantly higher VSC.
@@ B-C Stop Criterion Analysis
    1-NBV+MASCVP (Ours)     13.01 ± 1.46 ⋆⋆⋆98.41 ± 2.67 3.07 ± 0.24
TABLE XVI: Evaluation of random methods with a fixed number of views and our stop criterion: Each value is the mean value and its standard deviation, computed over 3,000 tests. RV, VSC, and MC stand for the number of required views, visible surface coverage, and movement cost, respectively. ⋆⋆⋆ indicates a consistent significance at $p<0.001$ between our method and all random methods (i.e., several paired tests), as determined by both paired t-test and the Wilcoxon signed-rank test. As can be seen, our 1-NBV+MASCVP achieves a significantly higher VSC against random methods, indicating our effective stop criterion.